A 20 L container is filled with helium and the pressure is 148 atm and the temperature is 318 K. What volume can be filled in liters when the temperature is 25° C and the atmospheric pressure is 760 torr?
1
Expert's answer
2019-03-15T02:35:38-0400

p1*V1/T1 = p2*V2/T2

p1 = 148 atm = 15000000 Pa

p2 = 760 torr = 101325 Pa


V2 = (p1 * V1 * T2) / (p2 * T1) = (15000000 * 20 * 318) / (101325 * 298) = 3160 L

Answer = 3160 L